OpenStudy (anonymous):

Let f(x) = −4x + 7 and g(x) = 10x − 6. Find f(g(x)).

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

so graph this and you will get your first awser y=−4x+7

OpenStudy (anonymous):

okay

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it does not say to graph, does it?

OpenStudy (anonymous):

and then hears how you do the 2nd part of your problem Let's solve for g. gx=10x−6 Step 1: Divide both sides by x. gxx=10x−6x g=10x−6x Answer: g=10x−6x

OpenStudy (misty1212):

lol oooooooh nooooooooo!

OpenStudy (anonymous):

y what am i doing wrong

OpenStudy (misty1212):

solve for \(g\)??!! darling, \(g\) is not a variable \(g(x)\) is a function it is \[g(x)=10x-6\]

OpenStudy (anonymous):

im lost

OpenStudy (misty1212):

\[f(x)=-4x+7\\ g(x)=10x-6\] so \[f(g(x))=f(10x-6)\] is the first step

OpenStudy (misty1212):

then since \[\huge f(\clubsuit)=-4\clubsuit +7\] you have \[\huge f(10x-6)=-4(10x-6)+7\]

OpenStudy (misty1212):

a little algebra will clean that up nicely

OpenStudy (anonymous):

would i just solve it like a regular two variable equation or how do i do it with f